2. A paper in the journal ‘Current Biology’ tells of the some jellyfish-like animals that attack their prey by launching stinging cells in one of the animal kingdom’s fastest movements. High-speed photography showed the cells were accelerated from rest for 700. ns at 5.30 X 10^7 m/s^2. Calculate (a) the maximum speed reached by the cells and (b) the distance traveled during the acceleration.
1
Expert's answer
2018-08-07T15:43:08-0400
(a) The maximum speed reached by the cells:
v=at=(5.30∙ 〖10〗^7 )(700∙〖10〗^(-9) )=37.1 m/s
(b) The distance traveled during the acceleration:
s=(at^2)/2=1/2 (5.30∙ 〖10〗^7 ) (700∙〖10〗^(-9) )^2=1.30 ∙〖10〗^(-5) m.